One side of a square of area 36 sq. cm. is coloured yellow. Find the length of the part which is

coloured yellow.


1
Expert's answer
2021-11-10T10:13:32-0500

Let "x=" the length of a side of the square, "x>0." Then the area "A" of a square will be


"A=x^2"

Given "A=36". Substitute


"x^2=36, x>0"

Hence "x=\sqrt{36}=6(cm)."

Therefore, the length of the part which is coloured yellow is 6 cm.